OpenStudy (elise_a18):

he regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. : 69 = x + one-half(x - 6) What is the regular price of a child's ticket? $38 $42 $48 $66

OpenStudy (elise_a18):

@sammixboo help pweeze?

sammixboo (sammixboo):

Well, a child's ticket price is 6 less than an adult ticket, so we need to solve for x, which is the original amount of the adult ticket first.

sammixboo (sammixboo):

Do you know how to solve for x? And it says one-half Is that 1/2?

OpenStudy (elise_a18):

yesh

sammixboo (sammixboo):

Okie dokie, so we have 69 = x + \(\dfrac{1}{2}\)f(x - 6) Do you know how to solve for x?

OpenStudy (elise_a18):

This problem is difficult for me though because It doesnt give me the price of the adult ticket so I have to do it backwards without a number to start with x(

sammixboo (sammixboo):

Oops, I left the F in there. Pretend that isn't there

OpenStudy (elise_a18):

I'm pretty sure I have to do something with the 69 reguarding the x

OpenStudy (elise_a18):

oops, regarding

sammixboo (sammixboo):

OK, so first for figuring out x, we must simplify the problem to make things easier 69 = x + \(\dfrac{1}{2}\)(x - 6) \(\downarrow\) 69 = \(\dfrac{3}{2}\)x - 3 Now let's solve for x

OpenStudy (elise_a18):

how did you get 3/2?

OpenStudy (elise_a18):

and where did the x go?